On the xy-coordinate plane, points A and B both lie on the circumference of a circle whose center is O, and the length of AB equals the circle's diameter. If the (x,y) coordinates of O are (2,1) and the (x,y) coordinates of B are (4,6), what are the (x,y) coordinates of A?
Answer by venugopalramana(3286)

On the xy-coordinate plane, points A and B both lie on the circumference of a circle whose center is O, and the length of AB equals the circle's diameter. If the (x,y) coordinates of O are (2,1) and the (x,y) coordinates of B are (4,6), what are the (x,y) coordinates of A?
O the centre of circle is mid point of diameter AB
hence if A is (x,y)
and B is (4,6)
O is (2,1)
so 2=(x+4)/2
x+4=4
x=0
and 1=(y+6)/2
y+6=2
y=-4
hence A is (0,-4)